OpenStudy (anonymous):

Debbie and Trisha are looking at the equation the square root of the quantity of 4 times x minus 6 equals square root of x . Debbie says that the solution is extraneous. Trisha says that the solution is non-extraneous. Is Debbie correct? Is Trisha correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (anonymous):

@radar

terenzreignz (terenzreignz):

Well, they can't possibly be both correct, since they pretty much contradict each other, so it has to be one or the other. \[\Large \sqrt{4x-6}=\sqrt x\] How did you go about solving this? :)

OpenStudy (anonymous):

idk thats the promblem

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Well, the radical signs (square roots) sort of screw things up, so why don't you square both sides of the equation to get rid of them? \[\Large \sqrt{p}^2=p\]

OpenStudy (anonymous):

so we sqare both sides?

terenzreignz (terenzreignz):

Yup.

terenzreignz (terenzreignz):

Then solve for x.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@radar

OpenStudy (radar):

What did you get when you squared both sides ?

OpenStudy (gorv):

ok lets write equation on left hand side square root of the quantity of 4 times x minus 6 first write without squareroot 4 times x= 4x and subtract 6 now .....so is 4x-6 now take square root \[\sqrt{4x-6}\]

OpenStudy (radar):